OpenStudy (anonymous):

10x^2 - 11x - 6 Factor

OpenStudy (anonymous):

(5x+2)(2x-3)

OpenStudy (anonymous):

how did u get that?

myininaya (myininaya):

This is in the form of ax^2+bx+c where a=10, b=-11, c=-6. Find factors of a*c that add up to be b. a*c=(10)(-6)=-60. What are factors of -60 that add up to be -11

OpenStudy (anonymous):

im not sure

myininaya (myininaya):

-15*4=-60 and -15+4=-11

myininaya (myininaya):

So replace -11x=-15x+4x

myininaya (myininaya):

so we have 10x^2-15x+4x-6

myininaya (myininaya):

Factor by grouping

myininaya (myininaya):

What do the first two terms have in common

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

this literally by observation. 10= 1*10= 2*5 factor of 6: 1*6=2*3 try mix and match the factors in form (ax+b)(cx+d) where a&c is a pair of factor 10 and b&d a pair of factor 6. and ad+bc=-11

OpenStudy (anonymous):

o okay

myininaya (myininaya):

You can also do trial factors like anfar if you like

OpenStudy (anonymous):

i also have the problem 125n^2 + 50n + 5

myininaya (myininaya):

ok so 5x(2x-3)+2(2x+3)

myininaya (myininaya):

(2x+3)(5x+2) which is what anfar got

myininaya (myininaya):

oops

myininaya (myininaya):

it was suppose to be 5x(2x-3)+2(2x-3)

myininaya (myininaya):

=(5x+2)(2x-3)

myininaya (myininaya):

now thats what anfar got

OpenStudy (anonymous):

for the 2nd problem: first, make the numbers smaller.note that 125,50 and 5 have common factors 5. so now u have: 5(25n^2+10n+1) then, u can try use the mix and match for the equation in the bracket.

OpenStudy (anonymous):

would it be 5(5n+ 2)....?

OpenStudy (anonymous):

no. it should be 5(5n+1)(5n+1)=5(5n+1)^2

OpenStudy (anonymous):

but then what happens to the 10?

OpenStudy (anonymous):

i put it this way: let say u have an equation in form Ax^2+Bx+C: u have to factor it in form of (ax+b)(cx+d) where: a*c=A , b*d=C and ad+bc=B

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

in the 2nd case, u considering 25n^2+10n+1 so, u hav A=25, B=10 and C=1 note that 1=1*1 or -1*-1 since B is positive, u take the positive ones. so, u now have b=d=1 thus, a*c=25 and a+c=10 now, 25=25*1=5*5. 25+1 not equal to 10. so u take a=c=5. therefore u have: 5(5n+1)(5n+1)=5(5n+1)^2 as the answer. u got it?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

good.. u r welcome.